Norman Corp. has prepared the sales budget.

The expected sales are as follows: $84,000 for March, $100,000 for April, $92,000 for May, $116,000 for June, and $98,000 for July.

Credit sales are 80% of total sales, and cash sales are remaining 20%. Collections of account receivables are 80% in the month of sale (and subject to 1% discount), 15% in the month following the sale. The remaining 5% is expected to be uncollectible.

(a) Prepare a schedule of cash collections for April through June.

(b) What is the ending accounts receivable balance on May 31?
Answer #1

Answer:

(a)

Month Sales Cash Sales (20% of Sales) Credit sales(Sales-Cash sales) Same month realisation(80% of Credit sales less 1% discount) Prevoius month credit sales realisation(15% of credit sales) Bad debts(5% of credit sales) Total Cash Received(C+E+F)
March 84,000 16,800 67,200 53,222.4 0 3,360 70,022.4
April 100,000 20,000 80,000 63,360 10,080 4,000 93,440
May 92,000 18,400 73,600 58,291.2 1,2000 3,680 88,691.2
June 116,000 23,200 92,800 73,497.6 11,040 4,640 107,737.6
July 98,000 19,600 78,400 62,092.8 13,920 3,920 95,612.8

(b)

Ending accounts receivables on may 31=

Receivables=$11,040 (15% of credit sales that will be collected in the next month)

Allowance for doubtful debts=$11,040 (Since it is mentioned that 5% are expected uncollectibles and actual bad debts have not been occured thus allowance=3360+4000+3680)

Net receivables=11040-11040=$0 (NIL)

If helpful, Thumbs UP please:)